let the quadratic equations x²+3x-k=0 has root a,b and x²+3x-10=0. has the root c,d such that modulus of the difference of the root of the first equations is equal to twice the modulus of the difference of the root of the second equations .

Grade:11

1 Answers

Shailendra Kumar Sharma
6 years ago
|a-b| =|c-d|
so (a-b)2 = (c-d)2
=> (a+b)2-4ab =(c+d)2-4cd
=> 9+4k=9+4*10
=> k=10 
so the value of k will be 10